# Large Language Model NLP Natural Language Processing

### v1

@@
Investigating the Robustness of Deductive Reasoning with Large Language Models
Fabian Hoppe1    Filip Ilievski1    Jan-Christoph Kalo2
1Vrije Universiteit Amsterdam
2Universiteit van Amsterdam
{f.hoppe, f.ilievski}@vu.nl, j.c.kalo@uva.nl
Abstract
Large Language Models have been shown to achieve impressive results for many reasoning-based Natural Language Processing (NLP) tasks, suggesting a degree of deductive reasoning capability. However, it remains unclear to which extent LLMs, in both informal and autoformalisation methods, are robust on logical deduction tasks. Moreover, while many LLM-based deduction methods have been proposed, there is a lack of a systematic study that analyses the impact of their design components. Addressing these two challenges, we propose the first study of the robustness of LLM-based deductive reasoning methods. We devise a framework with two families of perturbations: adversarial noise and counterfactual statements, which jointly generate seven perturbed datasets. We organize the landscape of LLM reasoners according to their reasoning format, formalisation syntax, and feedback for error recovery. The results show that adversarial noise affects autoformalisation, while counterfactual statements influence all approaches. Detailed feedback does not improve overall accuracy despite reducing syntax errors, pointing to the challenge of LLM-based methods to self-correct effectively.
@@ 1 Introduction
Deriving new knowledge from existing knowledge, as in deductive reasoning, is a key human cognitive skill necessary for various applications, including complex question-answering and decision-making. Deductive reasoning can be intensive for humans (e.g., taking a lot of time), require specific expertise (e.g., logicians), and lead to incorrect conclusions (e.g., due to biases). This promotes deductive reasoning over Natural Language (NL) as a key objective of human-centric AI. Automatic deduction engines aim to support humans by providing certifiable reasoning chains, avoiding invalid inferences, and accelerating the process. To provide effective support for deductive reasoning, AI must be able to formalise knowledge and rules provided in NL robustly.
Performing logical reasoning has received much interest in AI. In the early days, symbolic methods were aimed at transforming specific parts of language into logical statements Pereira (1982). Recently, LLMs have been shown to achieve impressive results for many reasoning-based NLP tasks, suggesting a degree of deductive reasoning capability Srivastava et al. (2023). In particular, generating informal reasoning chains via Chain-of-thought (CoT) prompting achieves good reasoning performance on many benchmarks Wei et al. (2022). Contrary to symbolic methods, LLMs can answer a deductive reasoning task without providing a formal intermediate reasoning chain. Nevertheless, these informal reasoning chains do not need to follow truth-preserving proof rules, thus leading to reasoning chains that are hard to verify. Recent work shows that many informal reasoning chains suffer from lack of faithfulness Ye and Durrett (2022); Tanneru et al. (2024).
@@
Addressing these challenges, autoformalisation approaches  Pan et al. (2023); Olausson et al. (2023) use LLMs to translate NL input into a logical form, and a deterministic symbolic solver to perform the deductive reasoning. Autoformalisation is thus a hybrid approach, which aims to provide a faithful and verifiable reasoning chain while leveraging the linguistic manipulation skills of LLMs. Autoformalisation faces two key challenges: First, since they translate rich NL into a limited grammar of symbols and operations, it is critical to leverage a syntax with an optimal tradeoff between translation accuracy and expressivity. Second, while autoformalisation chains provide an opportunity for syntactic and semantic validation and error analysis, designing an effective and efficient terror recovery mechanism is non-trivial. While prior autoformalisation systems leverage multiple syntaxes and error recovery mechanisms, no systematic study has investigated their impact on autoformalisation accuracy.
Meanwhile, a key requirement of LLM-based reasoning methods is their robustness to noise Ebrahimi et al. (2018) and out-of-distribution Hendrycks et al. (2021) inputs. Given the strong performance of LLMs across many domains and benchmarks Sarlin et al. (2020), dealing with noisy data in reasoning tasks has been considered more important Sourati et al. (2024). Most evaluations have focused on adversarial noise (e.g., lexical perturbations) Sarlin et al. (2020), while a recent study has also experimented with counterfactual statements Liu et al. (2023). While robustness evaluations for NLP tasks have yielded mixed results Wang and Zhao (2024); Liu et al. (2023), it remains unclear to which extent LLMs, in both informal and autoformalisation methods, are robust in logical deduction tasks.
We address these two challenges by investigating the robustness of LLM-based deductive reasoning methods. Our overall approach is summarized in Figure 1. Our study makes three contributions. First, following standard practices in evaluating robustness, we devise a robustness framework for logical deduction with two families of perturbations: adversarial noise, where the model needs to preserve its label in the face of added irrelevant information, and counterfactual perturbations, where a single alteration in the context flips the label of the question. The combinations of the perturbations produce seven variants from a given dataset. Second, we synthesize the landscape of existing LLM-based logical deduction methods into a methodological framework with three dimensions: reasoning format, grammar syntax, and error recovery mechanism. For each of these dimensions, we incorporate representative approaches in the literature. Third, We perform extensive experiments with seven LLMs on eight perturbed variants of a recent modular benchmark. Our findings provide nuanced insights into the robustness of LLM-based methods on logical deduction.
@@ 2.1 Methods for LLM-based Reasoning
Scaling up the size of LLMs enables strong performance in many NLP tasks by few-shot prompting Brown et al. (2020), which suggests inherent reasoning capabilities. CoT combines few-shot prompting with generating intermediate informal reasoning chains Wei et al. (2022). These informal reasoning skills motivated more elaborate prompting techniques, like Zero-Shot CoT Kojima et al. (2022) or self-consistency by generating multiple chains Wang et al. (2022b), as well as more complex structures than chains, such as Tree of Thoughts Yao et al. (2023) and Graph of Thoughts Besta et al. (2024). These methods use an LLM to generate intermediate steps and evaluate the output through self-refinement. Similarly, ProofWriter Tafjord et al. (2021) improves multi-hop reasoning by adding the intermediate results to the reasoning context Tafjord et al. (2021). A key benefit of informal reasoning chains is their flexibility, but this comes at the expense of guaranteeing faithfulness. Consequently, methods combining LLMs with formal reasoning have been suggested.
@@
Instead of informal reasoning, another way combines an LLM with a deterministic symbolic solver, e.g., a theorem prover. Here, the prover guarantees faithful and deterministic reasoning. This combination is known as autoformalisation. One of the first autoformalisation approaches using formal reasoning chains in combination with prompting is PAL Gao et al. (2023). The authors generate Python snippets alongside informal steps and generate the final response by executing the generated code snippets. Logic-LM Pan et al. (2023) prompts LLMs to generate multiple task-specific formalisations (logic programming, first-order logic, satisfiability modulo theories and constraint satisfaction), which are solved by dedicated solvers. They report higher robustness for longer reasoning chains compared to CoT reasoning. The extension Logic-LM++ Kirtania et al. (2024) tries to avoid new syntax errors by integrating a self-refinement mechanism. The LINC Olausson et al. (2023) approach uses the idea of self-consistency from CoT and generates multiple formalisations to avoid formalisation errors. Autoformalisation models achieve high accuracy for many deductive reasoning benchmarks, showing clear benefits for complex reasoning.
@@ 2.2 Robustness Evaluation of LLM Reasoning
The improvements of LLM-based reasoning have inspired the development of benchmarks investigating capabilities for deductive reasoning. The synthetic PrOntoQA Saparov and He (2022) dataset is built on modus ponens multi-hop reasoning and confirms reasoning capabilities for the largest models. However, they noted issues with proof planning and selecting correct proof steps for longer reasoning chains. The FOLIO Han et al. (2022) and AR-LSAT Zhong et al. (2022) benchmarks confirmed these errors for more complex reasoning and more naturalistic language. LogicBench Parmar et al. (2023) is a recent benchmark that systematically studies the performance of LLM-based reasoners across multiple inference rules (e.g., modus ponens), reporting a good model performance for predicate logic and first-order logic grammars. A vital challenge identified by these works is unfaithful reasoning, i.e., the intermediate steps are not used to infer the final result Ye and Durrett (2022); Lanham et al. (2023); Tanneru et al. (2024).
@@
A key requirement of LLM-based reasoning methods is their robustness to noise Ebrahimi et al. (2018) and out-of-distribution Hendrycks et al. (2021) inputs. Given the strong performance of LLMs across many domains and benchmarks Sarlin et al. (2020), dealing with noisy data in reasoning tasks has been considered more important Sourati et al. (2024), including adversarial and counterfactual perturbations. A variety of robustness tests have therefore been designed Wang et al. (2022a). These robustness benchmarks rely on the original problem’s perturbations through paraphrasing and distractions on character, word, and sentence levels Sourati et al. (2024); Sarlin et al. (2020). Many works try to generate adversarial examples to better understand the generalization capabilities of LLMs, which are shown to significantly decrease LLM performance Wang et al. (2022a). RUPbench is a recent robustness study on logical reasoning of LLMs Wang and Zhao (2024), covering many reasoning datasets and all three types of perturbations. Another category involves semantic changes in the input texts. Semantic changes can also be performed purely on a linguistic or logical level. Recent approaches use these logic-based perturbations Nakamura et al. (2023), which aligns with our robustness framework. Meanwhile, counterfactual studies of LLM robustness have been less common. One exception is RECALL Liu et al. (2023), a benchmark based on external knowledge bases, whose study reveals that LLMs are generally susceptible to external knowledge with counterfactual information and that simple mitigation strategies cannot significantly alleviate this challenge.
@@
We conduct the first investigation of robustness differences between formal and informal LLM-based logical deduction methods. For this purpose, we base our experiments on LogicBench Parmar et al. (2023), which systematically includes nine inference rules mapped to natural language situations. We develop seven additional variants of LogicBench incorporating adversarial noise and counterfactual statements, in line with prior work like Nakamura et al. (2023) that studies LLM robustness on other reasoning tasks.
@@ 3 Framework for Evaluating Robustness in Logical Deduction Tasks
Robustness types. Following prior work on investigating the robustness of LLMSarlin et al. (2020); Liu et al. (2023), we conceptualize two families of context perturbations. First, inspired by adversarial robustness studies Nakamura et al. (2023), we posit that a reasoning system must be robust to adversarial noise (distractions), i.e., including irrelevant information in the context should not alter its prediction. Noise can shift the focus of a text, making it more difficult for an LLM to capture the relevant context. Second, a reasoning system must be robust to counterfactual shifts in the context. Namely, the system must be faithful to the provided context without including biases from implicit world models Liu et al. (2023). If the context states that men are immortal, the reasoning must overwrite its belief of men as mortal. By introducing counterfactual perturbations contradicting common sense, we investigate whether LLM-based logical deduction methods use reasoning shortcuts or perform genuine reasoning. We focus our perturbations on the NL context $c$ rather than the question $q$ because it corresponds to background knowledge sources, which tend to vary significantly in real-world applications (e.g., compare reasoning based on a research article to reasoning over a list of facts). The perturbed task can be formalised as $f^{\prime}:(c^{\prime},q)\rightarrow\{true,false\}$.
@@ 3.1 Adversarial Noise
Formalisation. As monotonic reasoning, deductive reasoning must remain invariant to newly added irrelevant information, i.e., additional text that does not change the semantics of the premises used to derive the conclusion from the context. Let us consider a perturbed context that includes noisy information: $c^{\prime}=d_{1}\dots d_{k}c$, where each $d_{i}$ denotes a noisy sentence concatenated to $c$ for $k\in\{1,2,4\}$. The task function $f^{\prime}$ must resolve to the same output as the original function $f$ for its proof to remain valid. To avoid any distractions that might change the original semantics (e.g., by breaking inter-sentence co-references), we append distractions only to the beginning of the context.
Design of adversarial noise. We define three types of noise relevant to logical deduction tasks, which vary in their degree of referential content, formalisation complexity, and depth of logical reasoning. All noise sentences are sampled in a way that guarantees they do not impact the semantics of the context and the original proof. Figure 2 (bottom-left) shows examples of the three noise types.
@@
Encyclopedic (E) perturbations are NL sentences expressing factual information such as It is 21 km from Karimnagar, on the highway from Karimnagar to Peddapa. They express real-world information following pragmatic principles of language Grice (1975). Encyclopedic sentences are often difficult or even impossible to formalize in first-order logic. At the same time, encyclopedic facts are not connected by complex logical relations and lack the linguistic structure typical for reasoning contexts. In summary, they represent world information, have a high formalisation complexity, and have low logical reasoning depth.
@@
Logical (L) statements provide a typical structure of reasoning contexts and contain only knowledge that can be natively formalised, such as All dresses are clothes. Logical sentences include information about the world, albeit in a fictional form. The required formalisation usually requires more complex reasoning, like multi-hop inferences. Thus, logical perturbations introduce limited world information, have low formalisation complexity, and high reasoning depth.
@@
Tautological (T) perturbations are easily recognisable general statements, e.g., True is not false. They may include negations and a small number of disjunctions or conjunctions. As such, they contain no referential information and require simple reasoning. However, their formalisation is often difficult because many first-order logic syntaxes do not consider predefined truth constants, like $\bot$, $\top$. In summary, tautologies in NL contain no referential information, have high formalisation complexity, and have low reasoning depth.

### v2

# Investigating the Robustness of Deductive Reasoning with Large Language Models

@@
Abstract
Large Language Models have been shown to achieve impressive results for many reasoning-based Natural Language Processing (NLP) tasks, suggesting a degree of deductive reasoning capability. However, it remains unclear to which extent LLMs, in both informal and autoformalisation methods, are robust on logical deduction tasks. Moreover, while many LLM-based deduction methods have been proposed, a systematic study that analyses the impact of their design components is lacking. Addressing these two challenges, we propose the first study of the robustness of formal and informal LLM-based deductive reasoning methods. We devise a framework with two families of perturbations: adversarial noise and counterfactual statements, which jointly generate seven perturbed datasets. We organize the landscape of LLM reasoners according to their reasoning format, formalisation syntax, and feedback for error recovery. The results show that adversarial noise affects autoformalisation, while counterfactual statements influence all approaches. Detailed feedback does not improve overall accuracy despite reducing syntax errors, pointing to the challenge of LLM-based methods to self-correct effectively.
LLM
Large Language Model
NLP
Natural Language Processing
CoT
Chain-of-thought
QA
Question Answering
NL
Natural Language
FOL
First-Order Logic
AST
@@ Abstract
[…]
9195
@@ 1 Introduction
Deriving new knowledge from existing knowledge, as in deductive reasoning, is a key human cognitive skill necessary for various applications, including complex question-answering and decision-making. Deductive reasoning can be resource-intensive for humans (e.g., taking a lot of time), require specific expertise (e.g., logicians), and lead to incorrect conclusions (e.g., due to biases). This promotes automated deductive reasoning over Natural Language (NL) as a key objective of human-centric AI. Automatic deduction engines aim to support humans by providing certifiable reasoning chains, avoiding invalid inferences, and accelerating the process. To provide effective support for deductive reasoning, AI must be able to formalise knowledge and rules provided in NL robustly.
Performing logical reasoning has received much interest in AI. In the early days, symbolic methods were aimed at transforming specific parts of language into logical statements [20]. Recently, LLMs have been shown to achieve impressive results for many reasoning-based NLP tasks, suggesting a degree of deductive reasoning capability [25]. In particular, generating informal reasoning chains via Chain-of-thought (CoT) prompting achieves good performance on many benchmarks [32]. Contrary to symbolic methods, LLMs can answer a deductive reasoning task without providing a formal intermediate reasoning chain. Nevertheless, these informal reasoning chains do not need to follow truth-preserving proof rules, thus leading to chains that are hard to verify. Recent work shows that many informal reasoning chains suffer from a lack of faithfulness [34, 28].
@@
Addressing these challenges, autoformalisation approaches  [18, 17] use LLMs to translate NL input into a logical form, and a rigorous symbolic solver to perform the deductive reasoning. Autoformalisation is thus a hybrid approach, which aims to provide a faithful and verifiable reasoning chain while leveraging the linguistic manipulation skills of LLMs. Autoformalisation faces two key challenges: First, since they translate rich NL into a limited grammar of symbols and operations, it is critical to leverage a syntax with an optimal tradeoff between translation accuracy and expressivity. Second, while autoformalisation chains provide an opportunity for syntactic and semantic validation and error analysis, designing an effective and efficient error recovery mechanism is non-trivial. While prior autoformalisation systems leverage multiple syntaxes and error recovery mechanisms, no systematic study has investigated their impact on autoformalisation accuracy.
Meanwhile, a key requirement of LLM-based reasoning methods is their robustness to noise [3] and out-of-distribution [7] inputs. Given the strong performance of LLMs across many domains and benchmarks [23], dealing with noisy data in reasoning has been considered more important [24]. Most evaluations have focused on adversarial noise (e.g., lexical perturbations) [23], while a recent study has also experimented with counterfactual statements [15]. While robustness evaluations for NLP tasks have yielded mixed results [31, 15], it remains unclear to which extent LLMs, in both informal and autoformalisation methods, are robust in logical deductions.
We address these two challenges by investigating the robustness of LLM-based deductive reasoning methods with the goal to advance the understanding of LLM-based reasoning and shed light on methodological aspects to inform more complex, future reasoning systems. Our overall approach is summarised in Figure 1.
Our study makes three contributions. First, following standard practices in evaluating robustness, we devise a robustness framework for logical deduction with two families of perturbations: adversarial noise, where the model needs to preserve its label in the face of added irrelevant information, and counterfactual perturbations, where a single alteration in the context flips the label of the question. The combinations of the perturbations produce seven variants from a given dataset. The perturbed data is made available [8]. Second, we synthesise the landscape of existing LLM-based logical deduction methods into a methodological framework with three dimensions: reasoning format, grammar syntax, and error recovery mechanism. For each of these dimensions, we incorporate representative approaches in the literature and release the source code [8]. Third, we perform extensive experiments with seven LLMs on eight perturbed variants of a recent modular benchmark. Our findings provide nuanced insights into the robustness of LLM-based reasoning methods.
@@ 2.1 Methods for LLM-based Reasoning
Scaling up the size of LLMs enables strong performance in many NLP tasks by few-shot prompting [2], which suggests inherent reasoning capabilities. Chain-of-thought (CoT) combines few-shot prompting with generating intermediate informal reasoning chains [32]. These informal reasoning skills motivated more elaborate prompting techniques, like Zero-Shot CoT [13] or self-consistency by generating multiple chains [30], as well as more complex structures than chains, such as Tree of Thoughts [33] and Graph of Thoughts [1]. These methods use an LLM to generate intermediate steps and evaluate the output through self-refinement. Similarly, ProofWriter [27] improves multi-hop reasoning by adding the intermediate results to the reasoning context [27]. A key benefit of informal reasoning chains is their flexibility, but this comes at the expense of guaranteeing faithfulness. Consequently, methods combining LLMs with formal reasoning have been suggested.
@@
Instead of informal reasoning, another way combines an LLM with a deterministic symbolic solver, e.g., a theorem prover. Here, the prover guarantees faithful and deterministic reasoning. This combination is known as autoformalisation. One of the first autoformalisation approaches using formal reasoning chains in combination with prompting is PAL [4]. The authors generate Python snippets alongside informal steps and generate the final response by executing the generated code snippets. Logic-LM [18] prompts LLMs to generate multiple task-specific formalisations (logic programming, first-order logic, satisfiability modulo theories and constraint satisfaction), which are solved by dedicated solvers. They report higher robustness for longer reasoning chains compared to CoT reasoning. The extension Logic-LM++ [12] tries to avoid new syntax errors by integrating a self-refinement mechanism. The LINC [17] approach uses the idea of self-consistency from CoT and generates multiple formalisations to avoid formalisation errors. Autoformalisation models achieve high accuracy for many deductive reasoning benchmarks, showing clear benefits for complex reasoning.
@@ 2.2 Robustness Evaluation of LLM Reasoning
The improvements of LLM-based reasoning have inspired the development of benchmarks investigating capabilities for deductive reasoning. The synthetic PrOntoQA [22] dataset is built on modus ponens multi-hop reasoning and confirms reasoning capabilities for the largest models. However, they noted issues with proof planning and selecting the correct proof steps for longer reasoning chains. The FOLIO [6] and AR-LSAT [36] benchmarks confirmed these errors for more complex reasoning and more naturalistic language. LogicBench [19] is a recent benchmark that systematically studies the performance of LLM-based reasoners across multiple inference rules (e.g., modus ponens), reporting a good model performance for predicate logic and first-order logic. A vital challenge identified by these works is unfaithful reasoning, i.e., the intermediate steps are not used to infer the final result [34, 14, 28].
@@
A key requirement of LLM-based reasoning methods is their robustness to noise [3] and out-of-distribution [7] inputs. Given the strong performance of LLMs across many domains and benchmarks [23], dealing with noisy data in reasoning tasks has been considered more important [24], including adversarial and counterfactual perturbations. A variety of robustness tests have therefore been designed [29]. These robustness benchmarks rely on the original problem’s perturbations through paraphrasing and distractions on character, word, and sentence levels [24, 23]. Many works try to generate adversarial examples to better understand the generalisation capabilities of LLMs, which are shown to significantly decrease LLM performance [29]. RUPbench is a recent robustness study on logical reasoning of LLMs [31], covering many reasoning datasets and all three types of perturbations. Another category involves semantic changes in the input texts. Semantic changes can also be performed purely on a linguistic or logical level. Recent approaches use these logic-based perturbations [16], which align with our robustness framework. Meanwhile, counterfactual studies of LLM robustness have been less common. One exception is RECALL [15], a benchmark based on external knowledge bases, whose study reveals that LLMs are generally susceptible to external knowledge with counterfactual information and that simple mitigation strategies cannot significantly alleviate this challenge.
@@
We conduct the first investigation of robustness differences between formal and informal LLM-based logical deduction methods. For this purpose, we base our experiments on LogicBench [19], which systematically includes nine inference rules mapped to natural language situations. We develop seven additional variants of LogicBench incorporating adversarial noise and counterfactual statements, in line with prior work like [16] that studies LLM robustness on other reasoning tasks.
@@ 3 Framework for Evaluating Robustness in Logical Deduction Tasks
Robustness types. Following prior work on investigating the robustness of LLM[23, 15], we conceptualize two families of context perturbations. First, inspired by adversarial robustness studies [16], we posit that a reasoning system must be robust to adversarial noise (distractions), i.e., including irrelevant information in the context should not alter its prediction. Noise can shift the focus of a text, making it more difficult for an LLM to capture the relevant context. Second, a reasoning system must be robust to counterfactual shifts in the context. Namely, the system must be faithful to the provided context without including biases from implicit world models [15]. If the context states that men are immortal, the reasoning must overwrite its belief of men as mortal. By introducing counterfactual perturbations contradicting common sense, we investigate whether LLM-based logical deduction methods use reasoning shortcuts or perform genuine reasoning. We focus our perturbations on the NL context $c$ rather than the question $q$ because it corresponds to background knowledge sources, which tend to vary significantly in real-world applications (e.g., compare reasoning based on a research article to reasoning over a list of facts). The perturbed task can be formalised as $f^{\prime}:(c^{\prime},q)\rightarrow\{true,false\}$.
@@ 3.1 Adversarial Noise
Formalisation. As monotonic reasoning, deductive reasoning must remain invariant to newly added irrelevant information, i.e., additional text that does not change the semantics of the existing premises does not change the conclusion. Let us consider a perturbed context that includes noisy information: $c^{\prime}=d_{1}\dots d_{k}c$, where each $d_{i}$ denotes a noisy sentence concatenated to $c$ for $k\in\{1,2,4\}$. The task function $f^{\prime}$ must resolve to the same output as the original $f$ for its proof to remain valid. To avoid any distractions that might change the original semantics (e.g., by breaking inter-sentence co-references), we append distractions only to the beginning of the context.
Design of adversarial noise. We define three types of noise relevant to logical deduction tasks, which vary in their degree of referential content, formalisation complexity, and depth of logical reasoning. All noise sentences are sampled in a way that guarantees they do not impact the semantics of the context and the original proof. Figure 2 (top-right) shows examples of the three noise types.
@@
Encyclopedic (E) perturbations are NL sentences expressing factual information such as It is 21 km from Karimnagar, on the highway from Karimnagar to Peddapa. They express real-world information following pragmatic principles of language [5]. Encyclopedic sentences are often difficult or even impossible to formalise in first-order logic. At the same time, encyclopedic facts are not connected by complex logical relations and lack the linguistic structure typical for reasoning contexts. Consequently, this perturbations only has a low semantic similarity to the original context. In summary, they represent world information, have a high formalisation complexity, and have low logical reasoning depth.
@@
Logical (L) statements provide a typical structure of reasoning contexts and contain only knowledge that can be natively formalised, such as All dresses are clothes. Logical sentences include information about the world, albeit in a fictional form. As such, the semantic similarity is larger than encyclopedic perturbations due to the shared typical structure. The formalisation usually requires more complex reasoning, like multi-hop inferences. Thus, logical perturbations introduce limited world information, have low formalisation complexity, and high reasoning depth.
@@
Tautological (T) perturbations are easily recognisable correct statements, e.g., True is not false. They may include negations and one disjunction or conjunction. These perturbations provide the highest semantic similarity to the original context because they only contain the shared typical structure of the reasoning context without including information about the world. As such, they contain no referential information and require simple reasoning. However, their formalisation is often difficult because many first-order logic syntaxes do not consider predefined truth constants, like $\bot$, $\top$. In summary, tautologies in NL contain no referential information, have high formalisation complexity, and have low reasoning depth.